OpenStudy (anonymous):

if f(x)=8x and g(x)=2x+1, find(f o g)(x) A. 16x+8 B. 16x+1 C. 16x^2+8x D. 10x+1

OpenStudy (anonymous):

(f o g)(x) = f(g(x)) f(g(x)) = 8(2x + 1) = ?

OpenStudy (lgbasallote):

any further problems @SantanaG :)

OpenStudy (anonymous):

16x+1 so B @lgbasallote :)

OpenStudy (lgbasallote):

careful...8(2x+1)..you're gonna distribute 8 ;) 8(2x) + 8(1)

OpenStudy (anonymous):

oh yeahh so A lol

OpenStudy (lgbasallote):

correct-o! :D and lol give the medal to calcmathlete..he gave the key process =)) i just clarified the answer